# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ C.2 Geological Stratum
Tn and Tp are sedimentary rocks created during the Cenozoic era (shinseidai). Tp was created during a much earlier period, called the Paleogene period (ko daisanki), compared to Tn, which was made during the Neogene period (shin daisanki). Evidence indicates that Tp includes some coal layers created in the carboniferous ages (Nagao 1975). Figure C.4 shows a geological columnar section of the Iizuka coal mine, a representative coal mine in the Chikuho coalfield.
@@
Notes: The number of observations for 1931 and 1936 are $1,140$ and $1,364$, respectively.
@@
Note: This geological columnar section shows the strata around the Iizuka coal mine, including coal layers. The first to third columns include the names of layers created in the Paleogene period. The fourth to seventh columns examine the length of each layer, columnar, and legend. The black layers shown in the columnar of the sixth column are the coal layers: more than ten coal layers are included in the layers created in the Paleogene period.
@@ C.3 Dependent Variables
Notes: $\ddagger$ and $\dagger$ in Panels B–D indicate that the mean difference between treatment and control groups is statistically significant at the 1% and 5% levels, respectively. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine.
@@
Panel B reports the summary statistics for infant mortality, fetal death, and child mortality rates. The infant mortality rate is the number of infant deaths per $1,000$ live births. The fetal death rate is the number of fetal deaths per $1,000$ births. The child mortality rate is the number of deaths of children aged 1–4 per $1,000$ children aged 1–5. The variables with $\star$ have a set of censored observations; the results from the Tobit models are materially similar to the main results (not reported).
@@ C.3.3 Labor Supply
In Online Appendix D.4, I consider the labor force participation rate, the number of workers per $100$ people, to investigate the impacts on local labor supplies. To analyze structural shifts, I also use the employment share of workers in the mining, agricultural, manufacturing, commercial, and domestic sectors: these are calculated as the number of workers in each industry per $100$ workers. I consider these variables for both sexes to understand the potential gender bias in local labor supply and structural shifts. These labor statistics were obtained from the prefectural part of the 1930 Population Census. [^57] The prefectural part comprises $47$ reports (one for each prefecture), which I collected and digitized. For simplicity, I cite those reports as one document, Statistics Bureau of the Cabinet (1935a). The full population census is conducted every ten years in Japan; the 1930 census surveyed all labor statistics. Thus, similar statistics were unavailable in the 1935 Population Census. The 1940 Population Census was also disorganized because of the wartime regime. Panel C of Table C.4 presents the summary statistics of these variables by treatment status and census year. [^58] The employment share in the mining sector has a set of censored observations because mining is a localized economic activity. However, the estimates from the Tobit estimator (Tobin 1958) confirm that censoring does not affect the main results (not reported). Similar to the demographic outcomes, the mean differences are statistically significant for most variables, which implies the potential impact of mining extraction.
@@ C.4.1 Railway Stations
Source: Created by the author.
@@ C.4.2 Seaports
Notes: The blue circles indicate the location of seaports. The ports in Okinawa prefecture and a few ports in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
@@ C.4.3 Rivers
The location of rivers was obtained from the official shapefile provided by the Ministry of Land, Infrastructure, Transport, and Tourism (https://nlftp.mlit.go.jp/ksj/jpgis/datalist/KsjTmplt-W05.html, accessed August 21, 2018). The rivers used to match the mining dataset included those managed by the government or prefectures. As the rivers had not moved, the location information is time constant. Figure C.7a (C.7b) shows the distribution of the rivers matched to the 1931 (1936) base map.
@@
Source: Created by the author.
@@ C.4.4 Elevation
The grid cell data on elevation are taken from the official GIS database of the Ministry of Land, Infrastructure, Transport, and Tourism (https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-G04-d.html, accessed February 24, 2023). The average elevation in a 1 km square grid cell is calculated using the elevation measured in the 250 square grid cells in a third mesh-code initially assigned by the Ministry of Land, Infrastructure, Transport, and Tourism. The nearest neighbor grid cell is then linked to each municipality. The average distance to the nearest neighbor is $0.49$km, which is close enough to represent each municipality’s average elevation. Figure C.8 shows the spatial distribution of the elevation.
@@
Note: This figure illustrates the distribution of average elevation (in meters) in each municipality. The rivers in Okinawa prefecture and a few points in the islands far from the main island are not shown in the figures. The white circles indicate the location of coal mines in 1936. Source: Created by the author.
@@
Notes: This table shows the summary statistics for the control variables by type of mines. GSC mines indicate gold, silver, and copper mines. Distances to the nearest neighboring station, port, and river are in kilometers. Elevation is in meters.
@@ D.1 Attenuation due to Measurement Error
Consider a simplified projection of $y_{i}$ on $\textit{MineDeposit}_{i}$ as $y_{i}=\theta\textit{MineDeposit}_{i}+u_{i}$. When $\textit{MineDeposit}_{i}$ has a classical measurement error, as $\textit{MineDeposit}^{*}_{i}=\textit{MineDeposit}_{i}+u_{i}$, the observable model becomes $y_{i}=\theta^{*}\textit{MineDeposit}^{*}_{i}+\nu_{i}$. Consequently, I use $\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})=E[\textit{MineDeposit}^{*}_{i}\nu]=-\theta\text{var}(u_{i})\neq 0$ for $\theta^{*}=\theta+\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})/\text{var}(\textit{MineDeposit}^{*}_{i})$.
@@ D.2 Assumptions for the Identification
To simplify the discussion on the identification assumptions, I focus on the most straightforward structural equation considered in this study, a dummy endogenous variable model taking the form:
@@
All the variables are defined in Section 5.1. Note that I use the same Greek symbols as those used in equations 2 and 3 to simplify the expressions of the model, albeit those shall be different parameters to those considered in Section 5.1.
Five assumptions are made to identify the parameter of interest under this setting: the stable unit treatment value assumption (hereafter SUTVA), random assignments, exclusion restriction, nonzero average causal effect of the IV, and monotonicity (Angrist et al. 1996). Section 5.1 shows that the assumptions of the random assignments and exclusion restriction are plausible given the nature of the spatial distributions of mines and stratum. In this subsection, I then discuss the rest of the three assumptions.
First, the SUTVA argues that the potential outcomes for municipality $i$ in the system (20) are only related to its own treatment status. Under my empirical setting that utilizes the municipal-level dataset, this assumption is less likely to be violated because there are certain distances among different treated municipalities, even in the agglomerated area such as the Chikuhō coalfield. Moreover, to minimize the risk of SUTVA violation, I consider a certain threshold to define the treatment group. The results from the specifications show that the baseline definition of the treatment distance is plausible to model the potential spillover effects in the regression (Section D.6).
@@ D.3 Alternative Variance Estimator
To assess the potential influences of the local-scale spatial correlations, I use the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator (Arellano 1987). The finite-sample adjustment proposed by Hansen (2007) is implemented in the estimator. Table D.1 summarizes the results for the demographic outcomes, labor outcomes, and early-life health outcomes under the cluster-robust variance-covariance matrix estimator. All the results are materially similar to those of my baseline results in the main text.
@@
Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, and elevation.
@@ D.4 Gender-Biased Structure Shift: Evidence from the 1930 Census
In this section, I aim to evaluate how the coal mines influence the local industrial structure using labor supply statistics from the 1930 population census. Theory predicts the positive correlation between the resource extractions and local labor supply. Allcott and Keniston (2018) provides evidence of such a relationship for the oil and gas extraction industries in the post-World War II U.S. economy. My results show that the impacts of the coal mine on local labor supply depended on gender. While the local labor supply for male workers was not influenced by the coal extractions, that for female workers decreased with a meaningful magnitude. The rapid mechanization and institutional change observed in interwar Japan, which is not considered in the case of the postwar U.S. economy, may explain the gap in the findings. The labor-saving technological advances might had reduced the labor required in the mining sector and the labor regulations had further displaced women from the sector.
@@
Panel A of Table D.2 presents the results for the male workers. Column 1 shows the estimate for the male LFP is very close to zero and statistically insignificant. This indicates that coal mines did not influence the labor supply of male workers in the local economy. Columns 2–6 show the results for the employment share of male workers. Column 2 indicates that coal mines increased the mining sector’s employment share by approximately 34%, leading to a similar decline in the agricultural sector’s employment share (column 3). This suggests that the structural shift in male workers’ employment occurred mainly in the agricultural sector. In column 4, the estimate for the manufacturing sector is positive, suggesting a marginal spillover effect of mines on the manufacturing industry in the mining areas. However, this effect is not statistically significant. Compared with the case of the 1970s U.S., [^59] Black et al. (2005) found that the 1970s coal boom in the US increased employment and earnings with modest spillovers into the non-mining sectors. the rapid mechanization may not have created the time window for generating spillover effects in the case of interwar Japan. The estimates for the commercial and domestic sectors are moderately negative and close to zero (columns 5–6, respectively). This is consistent with male workers being less likely to work in these service sectors (Panel C of Table C.4).
@@
The results for female labor listed in Panel B show different responses. Column 1 shows a statistically significantly negative estimate, providing evidence that the coal mines decreased the female LFP. The estimate suggests that coal mines decreased the female LFP rate by approximately $19$%. This may be partly explained by the fact that the female relative wages in mining sector were relatively lower than those in the surrounding agrarian area (Panel B of Table 1). [^60] The clear gender difference in the labor force participation seems to be consistent with the wealth effect (“spending effect” in the Dutch-Disease literature) of mines: it may result in higher wage rates and lower overall non-resource GDP (Caselli and Coleman II 2001). Although it is difficult to analyze the impact of mines on wages, my results align with the findings on structural shifts presented in this subsection. Columns 2–6 show the results for the employment share of female workers. Column 2 indicates that coal mines increase the mining sector’s employment share by $24$% and decrease the agricultural sector’s share by the same degree (column 3). The estimated magnitude is smaller than that for male workers (Panel A). This makes sense because, on average, females were less likely to work in mines than males (Panel A of Table 1). While the manufacturing sector’s employment share decreases in the mining area (column 3), the estimates for the commercial and domestic sectors are moderately positive (columns 5–6, respectively). This seems to be in line with the increased relative demand for personal services from mining workers in mining areas (Caselli and Coleman II 2001).
@@ D.6 Testing the Validity of Threshold
Notes: This figure shows the results from the expanded regressions of equation 2 that includes the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (20–30] km bin is used as a reference group.
@@
I assess the validity of the threshold for the exposure variable (MineDeposit in equation 2) using my main outcome variables. In Figure D.1, I provide the results from the expanded regression of equation 2 that includes the five indicator variables for each 5 km bin from a mine. Figures D.1aD.1bD.1cD.1dD.1e, and D.1f show the results for the local population, marital fertility rate, female mortality rate, infant mortality rate, sex ratio, and average household size, respectively. Note that these estimates obtained from the OLS estimator are systematically attenuated toward zero, meaning that the estimates for the dependent variables with relatively moderate effects based on the IV estimator, such as fertility and mortality, are flatter. In addition, the estimates shown in these figures are the relative values to the reference group set, meaning that the statistical significance does not provide a meaningful interpretation of these figures. Therefore, the important point herein is that for the dependent variables with clear effects in the main text, the shape of the estimates shows a clear decreasing trend concerning the distance from mine. The population and sex ratio are particularly useful for this purpose. Overall, the estimates are larger in the absolute sense within $5$ km, whereas those for the areas outside the 5 km bin are close to zero and statistically insignificant. This evidence provides a basis for defining the exposure variable that uses $5$ km as the threshold.
@@
A 10-15 km bin for the infant mortality rate in 1938 shown in Figure D.1d shows a relatively clear positive estimate. To detect the region causing the hike, I split the entire sample into two subsamples: the municipalities belonging to Fukuoka prefecture and the others. The former might be an influential prefecture in the entire sample because it includes a number of mines in the Chikuhō coalfield. Figure D.2 shows the estimated coefficients for each subsample, confirming that the hike originated from the Fukuoka prefecture variations. Next, Figure D.3a illustrates the distribution of the bins in Fukuoka, and Figure D.3b clips the 10-15 km bin in the map. From a comparison of the distribution of the bins and the IMRs shown in Figure D.3c, it appears that some municipalities in the 10-15 km bin suffered a relatively higher risk of infant deaths, which is consistent with the tendency in Figure D.2. Note that, given a point of origin (i.e., a mine), a point on the concentric circle is a function of the distance. Therefore, the hike in the 10-15 km bin is considered to be a random event, given that the location of mines is at least conditionally independent as described in Section 5. Another evidence supporting this randomness is that such a hike in the 10-15km bin is not observed in the results for the 1933 sample (see Figure D.1d). If there were systematic factors, both 1933 and 1938 samples likely suffered the more significant risks of infant deaths in this bin. Similarly, the fact that only a 10-15 bin in Fukuoka prefecture suffers such a hike supports this argument. Overall, the hike is negligible in the sense that any systematic endogenous mechanism in the regression does not generate it.
@@
To summarize, the hike observed in the 10-15 km bin shall be randomly occurred, but not be led by any systematic factors that relate to the distribution of mines. [^61] Note that the models considered herein still allow the existence of unobservables that relate only to the IMR such as the social norms for the child-rearing because these socioeconomic factors unrelated to the distribution of mines. Although a large typhoon hit Miyazaki and Kagoshima prefectures in 1938, which might have increased the mortality rates, both prefectures are not included in my dataset. Given my empirical setting, this also makes sense because using many bins reduces the available observations in each bin, making the moment estimator vulnerable to this kind of random shock. [^62] As shown in Figure D.5, the infant mortality rates in Japan had moderate fluctuations during the 1930s, and the IMR in 1938 had slightly higher values than the surrounding years. This may further make the cross-sectional variation in the mortality rates greater.
@@
Notes: This figure shows the results from the expanded regressions of equation 2 that include the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (20–30] km bin is used as a reference group.

### v14

@@ C.2 Geological Stratum
Tn and Tp are sedimentary rocks created during the Cenozoic era (shinseidai). Tp was created during a much earlier period, called the Paleogene period (ko daisanki), compared to Tn, which was created during the Neogene period (shin daisanki). Evidence indicates that Tp includes coal layers created in the carboniferous ages (Nagao 1975; Tanaka 1984). Figure C.4 shows a geological columnar section of the Iizuka coal mine, a representative coal mine in the Chikuhō coalfield. This figure provides evidence that many coal layers are included in the layers created in the Paleogene period.
@@
Notes: This table summarizes the results of the linear probability models for the $19$ types of strata. The number of observations for 1931 and 1936 are $1,140$ and $1,364$, respectively.
@@
Note: This geological columnar section shows the strata around the Iizuka coal mine in Chikuhō coalfield in Kyūshū region. The first to third columns include the names of layers created in the Paleogene period. The fourth to seventh columns examine the length of each layer, columnar, and legend. The black layers shown in the columnar of the sixth column are the coal layers: more than ten coal layers are included in the layers created in the Paleogene period.
@@ C.3 Dependent Variables
Notes: $\ddagger$ and $\dagger$ in Panels A–C indicate that the mean difference between treatment and control groups is statistically significant at the 1% and 5% levels, respectively. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine.
@@
Panel B reports the summary statistics for infant mortality, fetal death, and child mortality rates. The infant mortality rate is the number of infant deaths per $1,000$ live births. The fetal death rate is the number of fetal deaths per $1,000$ births. The child mortality rate is the number of deaths of children aged 1–4 per $1,000$ children aged 1–5. The variables with $\star$ have a set of censored observations; the results from the Tobit models are materially similar to the main results.
@@ C.3.3 Labor Supply
In Online Appendix D.4, the labor force participation rate, defined as the number of workers per $100$ people, is considered to investigate the impacts on local labor supplies. To analyze structural shifts, the employment share of workers in the mining, agricultural, manufacturing, commercial, and domestic sectors are used. These are calculated as the number of workers in each industry per $100$ workers. These variables are considered for both sexes to understand the potential gender bias in the shifts. These labor statistics were obtained from the prefectural part of the 1930 Population Census. [^54] The prefectural part comprises $47$ reports (one for each prefecture), which I collected and digitized. For simplicity, those reports are cited as one document, Statistics Bureau of the Cabinet (1935a). The full population census is conducted every ten years in Japan. The 1930 census surveyed all labor statistics. Thus, similar statistics were unavailable in the 1935 Population Census. The 1940 Population Census was also disorganized because of the Second World War. Panel C of Table C.4 presents the summary statistics of these variables by treatment status and census year. [^55] The employment share in the mining sector has a set of censored observations because mining is a localized economic activity. As a robustness check, it was confirmed that the estimates from the Tobit estimator Tobin (1958) are similar to the main results (not reported). Similar to the demographic outcomes, the mean differences are statistically significant for most variables.
@@ C.4.1 Railway Stations
Source: Created by the author. The location data on the railway stations are from the Ministry of Land, Infrastructure, Transport and Tourism (website).
@@ C.4.2 Seaports
Notes: The blue circles indicate the location of the seaports. The ports in Okinawa prefecture and a few ports in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author. The location data on the seaports are from the Ministry of Land, Infrastructure, Transport and Tourism (website).
@@ C.4.3 Rivers
The location of rivers was obtained from the official shapefile provided by the Ministry of Land, Infrastructure, Transport, and Tourism (https://nlftp.mlit.go.jp/ksj/jpgis/datalist/KsjTmplt-W05.html, accessed August 21, 2018). The rivers used to match the mining dataset included those managed either by the government or prefectures. As the rivers had not moved, the location information is time constant. Figure C.7a (C.7b) shows the distribution of the rivers matched to the 1931 (1936) base map.
@@
Source: Created by the author. The location data of rivers are from the Ministry of Land, Infrastructure, Transport, and Tourism (website).
@@ C.4.4 Elevation
The grid cell data on elevation are taken from the official GIS database of the Ministry of Land, Infrastructure, Transport, and Tourism (https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-G04-d.html, accessed February 24, 2023). The average elevation in a 1 km square grid cell is calculated using the elevation measured in the 250 square grid cells in a third mesh-code initially assigned by the Ministry of Land, Infrastructure, Transport, and Tourism. The nearest neighbor grid cell is then linked to each municipality. The average distance to the nearest neighbor is $0.49$km, which is sufficiently close to represent each municipality’s average elevation. Figure C.8 shows the spatial distribution of the elevation.
@@
Note: This figure illustrates the distribution of average elevation (in meters) in each municipality. The rivers in Okinawa prefecture and a few points in the islands far from the main island are not shown in the figures. The white circles indicate the location of coal mines in 1936. Source: Created by the author. The grid cell data on elevation are from the Ministry of Land, Infrastructure, Transport, and Tourism (website).
@@
Notes: This table shows the summary statistics for the control variables by type of mines. GSC mines indicate gold, silver, and copper mines. Distances to the nearest station, port, and river are in kilometers. Elevation is in meters.
@@ D.1 Attenuation due to Measurement Error
Consider a simplified projection of $y_{i}$ on $\textit{MineDeposit}_{i}$ as $y_{i}=\theta\textit{MineDeposit}_{i}+\varepsilon_{i}$. When $\textit{MineDeposit}_{i}$ has a classical measurement error, $\textit{MineDeposit}^{*}_{i}=\textit{MineDeposit}_{i}+u_{i}$, where $u_{i}$ satisfies mean zero property. As explained in the main text, $u_{i}$ can be regarded as the unobservable changes in the assignments between the measured years of outcome and treatment variables. Given the random nature of the location of mines (Section 5.1), the measurement error ($u_{i}$) is likely to be uncorrelated with the outcome, true treatment variable, and error in the true model. Under this setting, the observable model becomes $y_{i}=\theta^{*}\textit{MineDeposit}^{*}_{i}+\nu_{i}$, where $\nu_{i}=\varepsilon_{i}-\theta u_{i}$. Consequently, I use $\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})=E[\textit{MineDeposit}^{*}_{i}\nu_{i}]=-\theta\text{var}(u_{i})\neq 0$ for $\theta^{*}=\theta+\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})/\text{var}(\textit{MineDeposit}^{*}_{i})$.
@@ D.2 Assumptions for the Identification
To simplify the discussion on the identification assumptions, I focus on the most straightforward structural equation, a dummy endogenous variable model taking the form:
@@
All the variables are defined in Section 5.1. Note that the same Greek symbols are used as those used in equations 2 and 3 to simplify the expressions of the model, albeit those shall be different parameters to those considered in Section 5.1.
Five assumptions are made to identify the parameter of interest under this setting: the stable unit treatment value assumption (hereafter SUTVA), random assignments, exclusion restriction, nonzero average causal effect of the IV, and monotonicity (Angrist et al. 1996). Section 5.1 shows that the assumptions of the random assignments and exclusion restriction are plausible given the nature of the spatial distributions of mines and stratum. This subsection then discusses the remaining three assumptions.
First, the SUTVA argues that the potential outcomes for municipality $i$ in the system (20) are only related to its own treatment status. Under my empirical setting that utilizes the municipal-level dataset, this assumption is less likely to be violated. This is because there are certain distances among different treated municipalities, even in the agglomerated area such as the Chikuhō coalfield. Moreover, to minimize the risk of SUTVA violation, a defined threshold is considered to define the treatment group. The results demonstrate that the baseline definition of the treatment distance is plausible to model the potential spillover effects in the regression (Online Appendix D.6).
@@ D.3 Alternative Variance Estimator
To assess the potential influences of the local-scale spatial correlations, the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator are used (Arellano 1987). The finite-sample adjustment proposed by Hansen (2007) is implemented in the estimator. Table D.1 summarizes the results based on the cluster-robust variance-covariance matrix estimator. All the results are materially similar to those of the baseline results in the main text.
@@
Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant mortality, fetal death, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, and elevation.
@@ D.4 Gender-Biased Structure Shift: Evidence from the 1930 Census
In this section, I aim to evaluate how the coal mines influence the local industrial structure using labor supply statistics from the 1930 population census. Theory predicts the positive correlation between the resource extractions and local labor supply. Allcott and Keniston (2018) provides evidence of such a relationship for the oil and gas extraction industries in the post-World War II United States (U.S.) economy. My results show that the impacts of the coal mine on local labor supply depended on gender. While the local labor supply for male workers was not influenced by the coal extractions, that for female workers decreased with a meaningful magnitude. The rapid mechanization and institutional change observed in interwar Japan, which is not considered in the case of the postwar U.S. economy, may explain the gap in the findings. The labor-saving technological advances might had reduced the labor required in the mining sector and the labor regulations had further displaced women from the sector.
@@
Panel A of Table D.2 presents the results for the male workers. Column 1 shows that the estimate for the male LFP is very close to zero and statistically insignificant. This indicates that coal mines did not influence the labor supply of male workers in the local economy. Columns 2–6 show the results for the employment share of male workers. Column 2 indicates that coal mines increased the mining sector’s employment share by approximately 34%, leading to a similar decline in the agricultural sector’s employment share (column 3). This suggests that the structural shift in male workers’ employment occurred mainly in the agricultural sector. In column 4, the estimate for the manufacturing sector is positive, suggesting a marginal spillover effect of mines on the manufacturing industry in the mining areas. However, this effect is not statistically significant. Compared with the case of the 1970s U.S., [^56] Black et al. (2005) found that the 1970s coal boom in the U.S. increased employment and earnings with modest spillovers into the non-mining sectors. the rapid mechanization may not have created the time window for generating spillover effects in the case of interwar Japan. The estimates for the commercial and domestic sectors are moderately negative and close to zero (columns 5–6, respectively). This is consistent with male workers being less likely to work in these service sectors (Panel C of Table C.4).
@@
The results for female labor listed in Panel B demonstrate different responses. Column 1 shows a statistically significantly negative estimate, providing evidence that the coal mines decreased the female LFP. The estimate suggests that coal mines decreased the female LFP rate by approximately $16$%. This may be partly explained by the fact that the gender wage gaps in mining sector were relatively lower than those in the surrounding agrarian area (Panel B of Table 1). [^57] The clear gender difference in the labor force participation appears to be consistent with the wealth effect (“spending effect” in the Dutch-Disease literature) of mines. It may result in higher wage rates and lower overall non-resource GDP (Caselli and Coleman II 2001). Although it is difficult to analyze the impact of mines on wages, the results align with the findings on structural shifts presented in this subsection. Columns 2–6 show the results for the employment share of female workers. Column 2 indicates that coal mines increase the mining sector’s employment share by $24$% and decrease the agricultural sector’s share by the same degree (column 3). The estimated magnitude is smaller than that for male workers (Panel A). This is logical because, on average, females were less likely to work in mines than males (Panel A of Table 1). While the manufacturing sector’s employment share decreases in the mining area (column 3), the estimates for the commercial and domestic sectors are moderately positive (columns 5–6, respectively). This appears to be in line with the increased relative demand for personal services from mining workers in mining areas (Caselli and Coleman II 2001).
@@ D.6 Testing the Validity of Threshold
Notes: This figure shows the results from the expanded regressions of equation 2 that includes the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (25–30] km bin is used as a reference group.
@@
The validity of the threshold for the exposure variable (MineDeposit in equation 2) is assessed using the main outcome variables. Figure D.1 provides the results from the expanded regression of equation 2 that includes the five indicator variables for each 5 km bin from a mine. Figures D.1aD.1bD.1cD.1dD.1e, and D.1f show the results for the local population, marital fertility rate, female mortality rate, infant mortality rate, sex ratio, and average household size, respectively. Note that these estimates obtained from the OLS estimator are systematically attenuated toward zero. This means that the estimates for the dependent variables with relatively moderate effects based on the IV estimator, such as fertility and mortality, are flatter. In addition, the estimates shown in these figures are the relative values to the reference group, meaning that the statistical significance does not provide a meaningful interpretation of these figures. Therefore, the important point herein is that the shape of the estimates for the dependent variables that captured clear effects in Section 5 does show a clear decreasing trend concerning the distance from mine. Therefore, the results for the population (Figure D.1a) and sex ratio (Figure D.1e) are useful for the purpose of this test. [^58] In other words, the other variables, such as fertility and mortality, are not useful herein because the main results do not have systematic effects. For example, the estimate for 1935 marital fertility is statistically insignificant in the main text (Table 3). This means that the estimates outside the 5km threshold can be similar to the estimate for the 0-5km bin. Figure D.1b confirms this relationship. Overall, the estimates are larger in the absolute sense within $5$ km, whereas those for the areas outside the 5 km bin are close to zero in Figures D.1a and D.1e. This evidence provides a basis for defining the exposure variable that uses $5$ km as the threshold.
@@
A 10-15 km bin for the IMR in 1938 shown in Figure D.1d shows a clear hike compared with the other estimates for the bins outside the 5km threshold. Since the main estimate for the 1938 IMR is not statistically significant (Table 5), this may be useful to analyze the reasons behind this hike. To detect the region causing the hike, the entire sample is stratified into two subsamples: the municipalities belonging to Fukuoka prefecture and the others. The former might be an influential prefecture in the entire sample because it includes a number of mines in the Chikuhō coalfield. Figure D.2 shows the estimated coefficients for each subsample, confirming that the hike originated from the Fukuoka prefecture variations. Next, Figure D.3a illustrates the distribution of the bins in Fukuoka, and Figure D.3b clips the 10-15 km bin in the map. From a comparison of the distribution of the bins and the IMRs shown in Figure D.3c, it appears that some municipalities in the 10-15 km bin suffered a relatively higher risk of infant deaths, which is consistent with the tendency in Figure D.2. Note that, given a point of origin (i.e., a mine), a point on the concentric circle is a function of the distance. Therefore, the hike in the 10-15 km bin is considered to be a random event, given that the location of mines is at least conditionally independent as described in Section 5. Another evidence supporting this randomness is that such a hike in the 10-15km bin is not observed in the results for the 1933 sample (see Figure D.1d). If there were systematic factors, both the 1933 and 1938 samples likely suffered the more significant risks of infant deaths in this bin. Similarly, the fact that only a 10-15 bin in Fukuoka prefecture suffers such a hike supports this argument. Overall, the hike is negligible because any systematic endogenous mechanism in the regression does not generate it.
@@
To summarize, the hike observed in the 10-15 km bin shall be randomly occurred, but may not be led by any systematic factors that relate to the distribution of mines. [^59] Note that the models considered herein still allow the existence of unobservables that relate only to the IMR. These include the social norms for the child-rearing because these socioeconomic factors unrelated to the distribution of mines. Although a large typhoon hit Miyazaki and Kagoshima prefectures in 1938, which might have increased the mortality rates, both prefectures are not included in the dataset. Given the empirical setting, this also makes sense because using many bins reduces the available observations in each bin, making the moment estimator vulnerable to this type of random shock. [^60] As shown in Figure D.5, the infant mortality rates in Japan had moderate fluctuations during the 1930s, and the IMR in 1938 had slightly higher values than the surrounding years. This may further make the cross-sectional variation in the mortality rates greater. As for the main analysis in Column (1) in Panel B of Table 5, this sort of hike in a specific bin should not be an influential event because the observations in the 10-15km bins, particularly of Fukuoka, become a small proportion of the control group.
@@
Notes: This figure shows the results from the expanded regressions of equation 2 that include the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (25–30] km bin is used as a reference group.